# Practical Rateless Set Reconciliation

### v2

@@ Abstract.
Set reconciliation, where two parties hold fixed-length bit strings and run a protocol to learn the strings they are missing from each other, is a fundamental task in many distributed systems. We present Rateless Invertible Bloom Lookup Tables (Rateless IBLT), the first set reconciliation protocol, to the best of our knowledge, that achieves low computation cost and near-optimal communication cost across a wide range of scenarios: set differences of one to millions, bit strings of a few bytes to megabytes, and workloads injected by potential adversaries. Rateless IBLT is based on a novel encoder that incrementally encodes the set difference into an infinite stream of coded symbols, resembling rateless error-correcting codes. We compare Rateless IBLT with state-of-the-art set reconciliation schemes and demonstrate significant improvements. Rateless IBLT achieves 33334×4$\times$4 × lower communication cost than non-rateless schemes with similar computation cost, and 22222000×2000$\times$2000 × lower computation cost than schemes with similar communication cost. We show the real-world benefits of Rateless IBLT by applying it to synchronize the state of the Ethereum blockchain, and demonstrate 5.6×5.6$\times$5.6 × lower end-to-end completion time and $\times4$ lower communication cost compared to the system used in production.
copyright: noneconference: ; journalyear: ;
@@ 1. Introduction
    Extensive experiments comparing Rateless IBLT with state-of-the-art solutions. Rateless IBLT achieves 33334×4$\times$4 × lower communication cost than regular IBLT (Goodrich and Mitzenmacher, 2011) and MET-IBLT (Lázaro and Matuz, 2023), two non-rateless schemes; and 22222000×2000$\times$2000 × lower computation cost than PinSketch (Dodis et al., 2004).
@@ 2. Motivation and Related Work
The performance of existing solutions varies depending on the codes they use. Characteristic Polynomial Interpolation (CPI) (Minsky et al., 2003) uses a variant of Reed-Solomon codes (Reed and Solomon, 1960), where coded symbols are evaluations of a polynomial uniquely constructed from $A$. CPI has a communication cost of $d\ell$, achieving the information-theoretic lower bound. However, its computation cost is $O(|A|d\ell)$ for Alice, and $O(|B|d\ell+d^{3}\ell^{4})$ for Bob. The latter was improved to $O(|B|d\ell+d^{2}\ell^{2})$ in PinSketch (Dodis et al., 2004; Wuille, [n. d.]) using BCH codes (Bose and Ray-Chaudhuri, 1960) that are easier to decode. Nevertheless, as we show in § 7.2, computation on both Alice and Bob quickly becomes intractable even at moderate $|A|$, $|B|$, and $d$, limiting its applicability. For example, Shrec (Han et al., 2020) attempted to use PinSketch to synchronize transactions in a high-throughput blockchain but found that its high computation complexity severely limits system throughput (Han et al., 2020, § 5.2).
@@ 3. Background
Limitations of IBLTs. IBLTs are not rateless. An IBLT with a particular set of parameters $m,k$ only works for a narrow range of difference size $d$. It quickly becomes inefficient to use it for more or fewer differences than parameterized for. In Appendix A, we show Theorems A.1 and A.2, which we summarize informally here. First, with high probability, Bob cannot recover any source symbol in $A\bigtriangleup B$ when $d>m$, making undersized IBLTs completely useless. On the other hand, we cannot simply default to a very large $m$ to accommodate a potentially large $d$. If $d$ turns out to be small, i.e., $d\ll m$, Alice still has to send almost the entire IBLT ($m$ coded symbols) for Bob to decode successfully, leading to high communication cost. Alice cannot dynamically enlarge $m$, either. Each source symbol is already uniformly mapped to $k$ out of $m$ coded symbols upon encoding. Increasing $m$ post hoc would require remapping the source symbols to the expanded space of coded symbols so that the mapping remains uniform. This requires Alice to rebuild and re-send the entire IBLT. Figs. 2(a)2(b) show an example.
@@ 4.1.2. Decodability
The following lemma shows that for this reason, the mapping probability $\rho(i)$ cannot decrease slower than $1/i^{1-\epsilon}$ for any positive $\epsilon$, i.e., almost inversely proportional to $i$. We defer the proof to Appendix C.
Lemma 4.1 ().
@@
The following lemma shows that for this reason, the mapping probability $\rho(i)$ cannot decrease faster than $1/i$. We defer the proof to Appendix C.
Lemma 4.2 ().
@@ 4.2. Realizing the Mapping Probability
Recall that for a random source symbol $s$, we want to make the probability that $s$ is mapped to the $i$-th coded symbol to be $\rho(i)$ in Eq. 1. A simple strawperson solution, for example, is to use a hash function that, given $s$, outputs a hash value uniformly distributed in $[0,1)$. We then compare the hash value to $\rho(i)$, and decide to map $s$ to the $i$-th coded symbol if the hash value is smaller. Given a random $s$, because its hash value distributes uniformly, the mapping happens with probability $\rho(i)$.
@@
We design an algorithm that maps each source symbol to the first $m$ coded symbols using only $O(\log m)$ computation. The strawperson solution is inefficient because we roll a dice (compare hash and $\rho(i)$) for every index $i$, even though we end up not mapping $s$ to the majority of them ($m-O(\log m)$ out of $m$), so reaching the next mapped index takes many dice rolls ($m/O(\log m)$ on average). Our key idea is to directly sample the distance (number of indices) to skip before reaching the next mapped index. We achieve it with constant cost per sample, so we can jump from one mapped index straight to the next in constant time.
@@
However, since there are $g$ (which can go to infinity) terms in $P_{g}$, it is still unclear how to sample $G$ in constant time. The key observation is that when we set the parameter $\alpha=0.5$ in $\rho(i)$, most of the terms in $P_{g}$ cancel out, and
$P_{g}=\frac{2(i+1)(i+2)}{(i+g)(i+g+1)(i+g+2)}.$
Let $C(x)$ be the cumulative mass function of $G$. It also has a remarkably simple form
$C(x)=\sum_{g=1}^{x}P_{g}=\frac{x(2i+x+3)}{(i+x+1)(i+x+2)}.$
We defer the step-by-step derivation of $P_{g}$ and $C(x)$ to Appendix B.
Let $C^{-1}(r)$ be the inverse of $C(x)$,
$C^{-1}(r)=\sqrt{\frac{(3+2i)^{2}-r}{4(1-r)}}-\frac{3+2i}{2}.$
In practice, we use the following approximation,
$C^{-1}(r)\approx(1.5+i)((1-r)^{-\frac{1}{2}}-1).$
To sample $G$, we sample $r\leftarrow[0,1)$ uniformly, and compute $g=\lceil C^{-1}(r)\rceil$. To make the algorithm deterministic, $r$ may come from a pseudorandom number generator seeded with the source symbol $s$. The algorithm outputs $i+g$ as the next index to which $s$ is mapped, updates $i\leftarrow i+g$, and is ready to produce another index. Because every source symbol is mapped to the first coded symbol (recall that $\rho(0)=1$), we start the recursion with $i=0$.
@@ 4.3. Resistance to Malicious Workload
In some distributed applications, a node may reconcile its state against multiple peers who are supposed to hold replicas of the system state. The common motivation is fault tolerance, where one node may go offline, forcing the recipient to receive updates from another; censorship resistance, where the recipient wants to make sure no peer hides information (e.g., social media posts from Alice, or her blockchain transactions); and efficiency, since updating simultaneously from multiple peers may be faster than choosing just one source.
@@
The literature on set reconciliation is aware of this issue, but typically does not specify the required properties from the hash function to mitigate it, or assumes strong properties such as random oracles (Mitzenmacher and Pagh, 2018), which have long outputs (say, 256 bits). It is sufficient, however, to use a keyed hash function with uniform and shorter outputs (say, 64 bits). This allows Alice and Bob to coordinate a secret key and use it to choose a hash function from the family of keyed hashes. Although with short hash images, an attacker can computationally enumerate enough symbols to find a collision for an item that Alice has, the attacker does not know the key, i.e., the hash function that Alice and Bob use, so she cannot target a collision to one of Alice’s symbols. This allows Rateless IBLT to minimize the size of a coded symbol and save bandwidth, particularly in applications where symbols are short and checksums account for much of the overhead. In practice, we use the SipHash (Aumasson and Bernstein, 2012) keyed hash function. A trade-off we make is that Alice has to compute the checksums separately for each key she uses, which increases her computation load. We believe this is a worthwhile trade-off as SipHash is very efficient, and we find in experiments (§ 7.2) that computing the hashes has negligible cost compared computing sums, which are still universal. Also, we expect using different keys only in applications where malicious workload is a concern.
@@ 5. Analysis
In this section, we use density evolution (Richardson and Urbanke, 2001; Luby et al., 1998) to analyze the behavior of the peeling decoder for symbols coded with Rateless IBLT. We mathematically prove that as the difference size $d$ goes to infinity, the overhead of Rateless IBLT converges to $1.35$; i.e., reconciling $d$ differences requires only the first $1.35d$ coded symbols. We then use Monte Carlo simulations to show that the convergence happens quickly, when $d$ is at the low hundreds.
Density evolution is a standard technique for analyzing the iterative decoding processes in error correcting codes (Richardson and Urbanke, 2001; Luby et al., 1998), and has been applied to IBLTs with simpler mappings between source and coded symbols (Lázaro and Matuz, 2021). Its high-level idea is to iteratively compute the probability that a random source symbol has not been recovered while simulating the peeling decoder statistically. If this probability keeps decreasing towards 00 as the peeling decoder runs for more iterations, then decoding will succeed with probability converging to $1$ (Luby et al., 1998, § 2). The following theorem states our main conclusion. In the interest of space, we defer its proof to Appendix C.
Theorem 5.1 ().
For a random set of $n$ source symbols, the probability that the peeling decoder successfully recovers the set using the first $\eta n$ coded symbols (as defined in § 4) tends to 1 as $n$ goes to infinity, provided that $\eta$ is any positive constant such that
@@
Theorem 5.1 implies that there exists a threshold $\eta^{*}$, which is the smallest $\eta$ that satisfies Eq. 2. Any $\eta>\eta^{*}$ also satisfies Eq. 2 because the left-hand side strictly decreases with respect to $\eta$. As long as Bob receives more than $\eta^{*}$ coded symbols per source symbol, he can decode with high probability. In other words, $\eta^{*}$ is the communication overhead of Rateless IBLT, i.e., the average number of coded symbols required to recover each source symbol. Recall that we set $\alpha=0.5$ in our final design as discussed in § 4.2. We solve for $\eta^{*}$ and get the following result.
Corollary 5.2 ().
@@
The average overhead of Rateless IBLT converges to $1.35$ as the difference size $d=|A\bigtriangleup B|$ goes to infinity.
In the appendix, we explore how $\alpha$ affects $\eta^{*}$ (see Fig. 14) and confirm that setting $\alpha=0.5$ is close to optimal.
@@
Theorem 5.1 and Corollary 5.2 from density evolution state the behavior of Rateless IBLT when the difference size $d$ goes to infinity. To understand the behavior when $d$ is finite, we run Monte Carlo simulations, and compare the results with the theorems. Fig. 4 shows the overhead at different $d$. It peaks at $1.72$ when $d=4$ and then converges to $1.35$, as predicted by Corollary 5.2. Convergence happens quickly: for all $d>128$, the overhead is less than $1.40$.
The fixed points of $q$ in Eq. 2 represent the expected fraction of source symbols that the peeling decoder fails to recover before stalling, as $d$ goes to infinity. Fig. 5 compares this result with simulations (we plot $1-q$, the fraction that the decoder can recover). The result matches closely to simulations. There is a sharp increase in the fraction of recovered source symbols towards the end, a behavior also seen in other codes that use the peeling decoder, such as LT codes (Luby, 2002).
@@
Figure 4. Overhead of Rateless IBLT at varying difference sizes $d$. We run 100 simulations for each data point. The shaded area shows the standard deviation. The dashed line shows $1.35$, the overhead predicted by density evolution.
@@
Figure 5. The fraction of recovered source symbols after receiving different number of coded symbols (normalized over the difference size $d$), as observed in simulations (1000 times), and as predicted by density evolution. Dashed line shows $1.35$, the overhead predicted by density evolution.

### v3

@@ Abstract.
Set reconciliation, where two parties hold fixed-length bit strings and run a protocol to learn the strings they are missing from each other, is a fundamental task in many distributed systems. We present Rateless Invertible Bloom Lookup Tables (Rateless IBLTs), the first set reconciliation protocol, to the best of our knowledge, that achieves low computation cost and near-optimal communication cost across a wide range of scenarios: set differences of one to millions, bit strings of a few bytes to megabytes, and workloads injected by potential adversaries. Rateless IBLT is based on a novel encoder that incrementally encodes the set difference into an infinite stream of coded symbols, resembling rateless error-correcting codes. We compare Rateless IBLT with state-of-the-art set reconciliation schemes and demonstrate significant improvements. Rateless IBLT achieves 33334×4$\times$4 × lower communication cost than non-rateless schemes with similar computation cost, and 22222000×2000$\times$2000 × lower computation cost than schemes with similar communication cost. We show the real-world benefits of Rateless IBLT by applying it to synchronize the state of the Ethereum blockchain, and demonstrate 5.6×5.6$\times$5.6 × lower end-to-end completion time and $\times4$ lower communication cost compared to the system used in production.
Set Reconciliation, Rateless Codes, Universal Codes, Data Synchronization, Randomized Algorithms
journalyear: 2024copyright: rightsretainedconference: ACM SIGCOMM 2024 Conference; August 4–8, 2024; Sydney, NSW, Australiabooktitle: ACM SIGCOMM 2024 Conference (ACM SIGCOMM ’24), August 4–8, 2024, Sydney, NSW, Australiadoi: 10.1145/3651890.3672219isbn: 979-8-4007-0614-1/24/08ccs: Networks Application layer protocolsccs: Networks Network algorithmsccs: Networks Network protocol designccs: Mathematics of computing Coding theoryccs: Computer systems organization Dependable and fault-tolerant systems and networks
@@ 1. Introduction
    Extensive experiments comparing Rateless IBLT with state-of-the-art solutions. Rateless IBLT achieves 33334×4$\times$4 × lower communication cost than regular IBLT (Goodrich and Mitzenmacher, 2011) and MET-IBLT (Lázaro and Matuz, 2023), two non-rateless schemes; and 22222000×2000$\times$2000 × lower computation cost than PinSketch (Dodis et al., 2008).
@@ 2. Motivation and Related Work
The performance of existing solutions varies depending on the codes they use. Characteristic Polynomial Interpolation (CPI) (Minsky et al., 2003) uses a variant of Reed-Solomon codes (Reed and Solomon, 1960), where coded symbols are evaluations of a polynomial uniquely constructed from $A$. CPI has a communication cost of $d\ell$, achieving the information-theoretic lower bound. However, its computation cost is $O(|A|d\ell)$ for Alice, and $O(|B|d\ell+d^{3}\ell^{4})$ for Bob. The latter was improved to $O(|B|d\ell+d^{2}\ell^{2})$ in PinSketch (Dodis et al., 2008; Wuille, 2018) using BCH codes (Bose and Ray-Chaudhuri, 1960) that are easier to decode. Nevertheless, as we show in § 7.2, computation on both Alice and Bob quickly becomes intractable even at moderate $|A|$, $|B|$, and $d$, limiting its applicability. For example, Shrec (Han et al., 2020) attempted to use PinSketch to synchronize transactions in a high-throughput blockchain but found that its high computation complexity severely limits system throughput (Han et al., 2020, § 5.2).
@@ 3. Background
Limitations of IBLTs. IBLTs are not rateless. An IBLT with a particular set of parameters $m,k$ only works for a narrow range of difference size $d$. It quickly becomes inefficient to use it for more or fewer differences than parameterized for. In § A, we show Theorems A.1 and A.2, which we summarize informally here. First, with high probability, Bob cannot recover any source symbol in $A\bigtriangleup B$ when $d>m$, making undersized IBLTs completely useless. On the other hand, we cannot simply default to a very large $m$ to accommodate a potentially large $d$. If $d$ turns out to be small, i.e., $d\ll m$, Alice still has to send almost the entire IBLT ($m$ coded symbols) for Bob to decode successfully, leading to high communication cost. Alice cannot dynamically enlarge $m$, either. Each source symbol is already uniformly mapped to $k$ out of $m$ coded symbols upon encoding. Increasing $m$ post hoc would require remapping the source symbols to the expanded space of coded symbols so that the mapping remains uniform. This requires Alice to rebuild and re-send the entire IBLT. Figs. 3(a)3(b) show an example.
@@ 4.1.2. Decodability
The following lemma shows that for this reason, the mapping probability $\rho(i)$ cannot decrease slower than $1/i^{1-\epsilon}$ for any positive $\epsilon$, i.e., almost inversely proportional to $i$. We defer the proof to § C.
Lemma 4.1.
@@
The following lemma shows that for this reason, the mapping probability $\rho(i)$ cannot decrease faster than $1/i$. We defer the proof to § C.
Lemma 4.2.
@@ 4.2. Realizing the Mapping Probability
Recall that for a random source symbol $s$, we want to make the probability that $s$ is mapped to the $i$-th coded symbol to be $\rho(i)$ in Eq. 1. A simple strawman solution, for example, is to use a hash function that, given $s$, outputs a hash value uniformly distributed in $[0,1)$. We then compare the hash value to $\rho(i)$, and decide to map $s$ to the $i$-th coded symbol if the hash value is smaller. Given a random $s$, because its hash value distributes uniformly, the mapping happens with probability $\rho(i)$.
@@
We design an algorithm that maps each source symbol to the first $m$ coded symbols using only $O(\log m)$ computation. The strawman solution is inefficient because we roll a dice (compare hash and $\rho(i)$) for every index $i$, even though we end up not mapping $s$ to the majority of them ($m-O(\log m)$ out of $m$), so reaching the next mapped index takes many dice rolls ($m/O(\log m)$ on average). Our key idea is to directly sample the distance (number of indices) to skip before reaching the next mapped index. We achieve it with constant cost per sample, so we can jump from one mapped index straight to the next in constant time.
@@
However, since there are $g$ (which can go to infinity) terms in $P_{g}$, it is still unclear how to sample $G$ in constant time. The key observation is that the cumulative mass function of $G$, denoted as $C(x)$, has a remarkably simple form. In particular,
(2) $C(x)=\sum_{g=1}^{x}P_{g}=1-\frac{\Gamma(i+1+\frac{1}{\alpha})\Gamma(x+i+1)}{\Gamma(i+1)\Gamma(x+i+1+\frac{1}{\alpha})}.$
We defer the step-by-step derivation to § B.
Let $C^{-1}(r)$ be the inverse of $C(x)$. The simple form of $C(x)$ allows us to compute $C^{-1}(r)$ easily, which we will soon explain. To sample $G$, we sample $r\leftarrow[0,1)$ uniformly, and compute $g=\lceil C^{-1}(r)\rceil$. To make the algorithm deterministic, $r$ may come from a pseudorandom number generator seeded with the source symbol $s$. The algorithm outputs $i+g$ as the next index to which $s$ is mapped, updates $i\leftarrow i+g$, and is ready to produce another index. Because every source symbol is mapped to the first coded symbol (recall that $\rho(0)=1$), we start the recursion with $i=0$.
Finally, we explain how to compute $C^{-1}(r)$. It is the most simple if we set the parameter $\alpha$ in $\rho(i)$ to $0.5$. Plugging $\alpha=0.5$ into Eq. 2, we get
$C(x)=\frac{x(2i+x+3)}{(i+x+1)(i+x+2)}.$
Its inverse is
$C^{-1}(r)$ $=\sqrt{\frac{(3+2i)^{2}-r}{4(1-r)}}-\frac{3+2i}{2}$
$\approx(1.5+i)((1-r)^{-\frac{1}{2}}-1).$
For a generic $\alpha$, we can use Stirling’s approximation (Robbins, 1955), and get
$C(x)\approx 1-\left(\frac{i+1}{x+i+1}\right)^{\frac{1}{\alpha}}.$
Consequently,
$C^{-1}(r)\approx(i+1)((1-r)^{-\alpha}-1).$
In our final design, we set $\alpha=0.5$. The main reason is that computing $C^{-1}(r)$ when $\alpha=0.5$ only requires computing square roots, while it otherwise involves raising $1-r$ to other non-integer powers. We observe that the latter is significantly slower on older CPUs. Meanwhile, as we will show in § 5, setting $\alpha=0.5$ results in negligible extra communication compared to the optimal setting.
@@ 4.3. Resistance to Malicious Workload
The literature on set reconciliation is aware of this issue, but typically does not specify the required properties from the hash function to mitigate it; most use hash functions with strong properties such as random oracles (Mitzenmacher and Pagh, 2018), which have long outputs (e.g., 256 bits). It is sufficient, however, to use a keyed hash function with uniform and shorter outputs (e.g., 64 bits). This allows Alice and Bob to coordinate a secret key and use it to choose a hash function from the family of keyed hashes. Although with short hashes, an attacker can computationally enumerate enough symbols to find a collision for an item that Alice has, the attacker does not know the key, i.e., the hash function that Alice and Bob use, so she cannot target a collision to one of Alice’s symbols. This allows Rateless IBLT to minimize the size of a coded symbol and save bandwidth, particularly in applications where symbols are short and checksums account for much of the overhead. In practice, we use the SipHash (Aumasson and Bernstein, 2012) keyed hash function. A trade-off we make is that Alice has to compute the checksums separately for each key she uses, which increases her computation load. We believe this is a worthwhile trade-off as SipHash is very efficient, and we find in experiments (§ 7.2) that computing the hashes has negligible cost compared to computing sums, which are still universal. Also, we expect using different keys only in applications where malicious workload is a concern.
@@ 5. Analysis
In this section, we use density evolution (Richardson and Urbanke, 2001; Luby et al., 1998) to analyze the behavior of the peeling decoder when decoding coded symbols in Rateless IBLTs. We mathematically prove that as the difference size $d$ goes to infinity, the overhead of Rateless IBLTs converges to $1.35$; i.e., reconciling $d$ differences requires only the first $1.35d$ coded symbols. We then use Monte Carlo simulations to show the behavior for finite $d$. In particular, we show that the overhead converges quickly, when $d$ is at the low hundreds.
Density evolution is a standard technique for analyzing the iterative decoding processes in error correcting codes based on sparse graphs (Richardson and Urbanke, 2001; Luby et al., 1998), and has been applied to IBLTs with simpler mappings between source and coded symbols (Lázaro and Matuz, 2021). Its high-level idea is to iteratively compute the probability that a random source symbol has not been recovered while simulating the peeling decoder statistically. If this probability keeps decreasing towards 00 as the peeling decoder runs for more iterations, then decoding will succeed with probability converging to $1$ (Luby et al., 1998, § 2). The following theorem states our main conclusion. We defer its proof to § C.
Theorem 5.1.
For a random set of $n$ source symbols, the probability that the peeling decoder successfully recovers the set using the first $\eta n$ coded symbols (as defined in § 4.1) tends to 1 as $n$ goes to infinity, provided that $\eta$ is any positive constant that satisfies
(3) $\forall q\in(0,1]:e^{\frac{1}{\alpha}\text{Ei}\left(-\frac{q}{\alpha\eta}\right)}<q.$
@@
Theorem 5.1 implies that for any choice of parameter $\alpha$, there exists a corresponding threshold $\eta^{*}$ which is the smallest $\eta$ that satisfies Eq. 3. Any $\eta>\eta^{*}$ also satisfies Eq. 3 because the left-hand side monotonically decreases with respect to $\eta$. (Intuitively, this must be true as a larger $\eta$ means more coded symbols, which should be strictly beneficial for decoding.) As long as Bob receives more than $\eta^{*}$ coded symbols per source symbol, he can decode with high probability. In other words, $\eta^{*}$ is the communication overhead of Rateless IBLTs, i.e., the average number of coded symbols required to recover each source symbol. $\eta^{*}$ is a function of $\alpha$. As discussed in § 4.2, we set $\alpha=0.5$ in our final design to simplify the process of generating mappings according to $\rho(i)$. We solve for $\eta^{*}$ when $\alpha=0.5$ and get the following result.
Corollary 5.2.
@@
The average overhead of Rateless IBLTs converges to $1.35$ as the difference size $d=|A\bigtriangleup B|$ goes to infinity.
5.1. Monte Carlo Simulations
@@
Theorem 5.1 and Corollary 5.2 from the density evolution analysis state the behavior of Rateless IBLTs when the difference size $d$ goes to infinity. To understand the behavior when $d$ is finite, we run Monte Carlo simulations, and compare the results with the theorems.
Fig. 4 shows the main results. It compares the overhead predicted by Theorem 5.1 and that observed in simulations. First, notice that as the difference size increases, simulation results converge to the analysis for all $\alpha$. How fast the results converge depends on $\alpha$. For all $\alpha\leq 0.55$, convergence happens quickly, and the overhead observed in simulations stays within $10\%$ of the analysis even for the smallest difference size we test. On the other hand, for $\alpha=0.95$, simulation results are still $12\%$ higher than the analysis at the largest difference size we test. Second, the figure shows that setting $\alpha=0.5$ is close to optimal for the communication overhead. Setting $\alpha=0.5$ results in $\eta^{*}=1.35$, while the optimal setting is $\alpha=0.64$ which results in $\eta^{*}=1.31$, a difference of only $3\%$.
Figure 4. Relationship between the communication overhead $\eta^{*}$ and the parameter $\alpha$ in $\rho(i)$. “DE” shows results from the density evolution analysis which assumes the difference size goes to infinity. Points show results from Monte Carlo simulations for various finite difference sizes. Each point is the average over 100 runs.
Next, we focus on $\alpha=0.5$, the parameter we choose for our final design. Fig. 5 shows the overhead as we vary the difference size $d$. It peaks at $1.72$ when $d=4$ and then converges to $1.35$ as predicted by Corollary 5.2. Convergence happens quickly: for all $d>128$, the overhead is less than $1.40$.
@@
Figure 5. Overhead of Rateless IBLTs at varying difference sizes $d$. We run 100 simulations for each data point and report the average. The shaded area shows the standard deviation. The dashed line shows $1.35$, the overhead predicted by density evolution.
The density evolution analysis also predicts how decoding progresses as the decoder receives more coded symbols. The fixed points of $q$ in Eq. 3 represent the expected fraction of source symbols that the peeling decoder fails to recover before stalling, as $d$ goes to infinity. Fig. 6 compares this result with simulations (we plot $1-q$, the fraction that the decoder can recover) and they match closely. There is a sharp increase in the fraction of recovered source symbols towards the end, a behavior also seen in other codes that use the peeling decoder, such as LT codes (Luby, 2002).
@@
Figure 6. The fraction of recovered source symbols after receiving different number of coded symbols (normalized over the difference size $d$), as observed in simulations (average of 1000 runs), and as predicted by density evolution. Dashed line shows $1.35$, the overhead predicted by density evolution.